Tuesday, December 16, 2008

Station Destroyed

This station was apparently destroyed by Hurricane Paloma tossing debris (or a boat) against the pylon during the first week in November, 2008, and is currently being considered for recommissioning by the Caribbean Community Climate Change Center based in Belize. To see old data collected for this site, please visit:

http://ecoforecast.coral.noaa.gov/index/0/DBJM1/station-home

For further information, please write to Dr. Kenrick R Leslie .

Sunday, July 13, 2008

Jamaica station back online

[note: copied from an email message sent on July 13th.]

Good news, yesterday the transmissions from the Jamaica station resumed. This was Saturday, July 12, at about 8am (Jamaica local time). Battery levels are extremely low but climbing. This means that the station was offline for 16 days, and took approximately two weeks to recover after the failed light sensor was removed.

Damage done: right now, none of the light sensors are reporting. I have no idea why this may be the case. Obviously there is no deep light sensor connected, so there's no surprise there. For the shallow and surface light sensors, they may have suffered some sort of irreversible damage when there was a short circuit in the station electronics. It's still possible they will revive, although it doesn't seem very likely to me.

The CTDs and the Vaisala WXT are reporting, although they've all now reached our usual instrument deployment lifetime and should be replaced.

Tuesday, July 01, 2008

DBJM1 power failure, Deep BIC implicated

[note: copied from an email message sent on July 1st.]

Many of you will have noticed that the Jamaica station is offline. I've examined the last few weeks of data and I've formed a hypothesis about the cause of failure, and I'm also including a timeline of the events leading up to the failure. The "executive summary" is this: I believe that a failure of the Deep BIC has in turn led to a station-wide power failure.

My recommendation is that DBML should send someone immediately to remove the Deep BIC. There is a chance, perhaps small, that this will bring the station back online. Standard instrument-replacement procedures should still be followed, i.e., the Deep BIC should be retrieved to a boat before disconnecting and the cable's end should be dummy-plugged and then secured to the pylon. [We may need to send them a female-style dummy plug if they don't have any on site.]

They should also try to inspect the entire length of the Deep BIC's cable from where it exits the pylon to its end, and look for places where it may have worn through. [If the problem is the cable, which in my view is less likely, then this requires disconnection from the brain unit and eventual cable replacement, both of which require a pylon-climb. Shoe might have some suggestions about what could be done with a worn cable at the waterline if a pylon-climber isn't immediately available.]

It would also be helpful if Peter could try to connect to the station via the radio unit and see if it responds. This could tell us whether the station is still running on enough power to log its data locally or not.

What do we know about the cause? Many instruments report supply voltages (the logger, the transmitter, the BICs, the WXT), and all of them show atypical power patterns beginning in early June. This was followed by the failure of the Deep BIC, a brief return to normal power patterns, and a final, dramatic drop in power levels last Wednesday and Thursday. The station has been completely offline since then.

What is the station's current status? The implications of that final power loss are troubling: if the station can't power the datalogger, then it isn't logging its data locally either. In this case only the CTDs will continue to run, and they may run out of battery power in a few months. On the other hand, it may just be that the station cannot supply the required voltage and current to power the transmitter during its transmission cycle. In this case, some local data collection and storage might still be going on.

What is wrong with the BIC? The most likely explanation is that there has been a bulkhead failure where the connector is attached. This has happened to the underwater BICs a few times before. If there has been ocean intrusion into the BIC's case, station power/ground wires may be effectively shorted within the instrument. If so, removing the BIC and dummy-plugging its cable may revive the station because it will open the short circuit and the station can recharge its depleted batteries with the solar panels.

What about the work done last week? The failure is part of a series of events that began on June 8th, so the events of last week are at most an aggravating factor. Also, there was a station cleaning in early June but a few days went by before the trouble began so that visit is not clearly implicated either. I haven't seen a blog entry from that cleaning so I'm not sure whether they noted anything out of the ordinary about the Deep BIC at that time.

Here is the failure timeline, with all times given in local Jamaica time (UTC-5). Times are approximate, accurate to within about an hour.

Tue Jun 03, 07am : groundtruth CT connected, presumably for cleaning
Wed Jun 04, 12pm : groundtruth CT disconnected
Sun Jun 08, 08pm : Deep BIC starts reporting unusually low voltages
Mon Jun 16, 08am : Deep BIC starts missing measurements
Tue Jun 17, 08pm : Deep BIC goes offline (except for one brief revival)
Sat Jun 21, 09am : Station power cycles return to normal patterns
Wed Jun 25, 06pm : Station begins final power-level plunge
Thu Jun 26, 11am : Station offline completely

regards,
Mike J+

Monday, April 28, 2008

Station Maintenance

15th April - calibration CTD hooked up @ 1045 hrs & removed @ 0900hrs 17th April

Full station cleaning carried out between 1430 & 1630 hrs
one of copper mesh covers on calib'n CTD exchanged for one from lower CTD unit which was growing a hole

Some way to be found to facilitate easy cleaning of area underneath the screen
on the CO2 unit ......Lobophora and Enteromorpha spp growing well under there

Peter G

Wednesday, February 27, 2008

Major Station Cleaning

1210 - 1350 major station cleaning carried out
Zinc anodes seemed to be decaying at faster rate than when station first installed.
4 Zincs on chains replaced
SAMI zinc will need to be replaced at next cleaning

Tuesday, February 12, 2008

Basic Station Cleaning

1110 - 1245hrs Basic station cleaning carried out

4 x zincs on chain replaced
2 x CTD zincs replaced
2 x Cu screen on CTDs replaced

Thursday, January 17, 2008

Major Station Cleaning

1225 - 1340 Major station cleaning carried out

Friday, November 30, 2007

Full Station Cleaning

1258 - 1425hrs.......full station cleaning carried out......biofouling still appeared minimal despite interval since last cleaning........when the calibration unit was first connected the readings form the data upload did not register and we had to go back out , disconnect and reconnect the unit before the system picked up and registered its presence......hence the delay in cleaning and the calibration unit was not removed from the system until the following day
Because of the slightly increased tilt to the unit, one of the rope legs (Se'n) became slightly slackened and was rubbing on a nearby coral head and suffered some (apparently minimal) abrasion damage. ......we removed the offending piece of coral and added more rope to the system to provide more support but reckon the tilting is the problem.

Thursday, October 25, 2007

Basic Cleaning of Station

basic cleaning 1015 - 1145 hrs

Biofouling appeared minimal

Sunday, September 16, 2007

Repairs To Pylon

Divers Jules Craynock and Derek Manzello inspected DBJM1 on Sept.11, 2007. Pin# 8 showed cracked substrate and concrete exposing 25% of the upper pin body. The T- pin was movable in bending toward the pylon, but not twisting. An approximate 5 degree listing of the pylon occurs, opposite pin 8 in a SE direction. All lashings and legs are in good condition and the installation is stable for normal conditions. Pin 1 showed slight upper substrate cracks but is sturdy. On Sept. 12, 2 qts. of 2-part u/w epoxy were applied to Pin 8 to rebuild the cracked substrate and the repair appears to have stopped the bending movement. The leg to pin 8 will have to be retensioned in the future and the behavior of the repaired pin under the tension needed to right the pylon is uncertain. No twisting of the pylon had occurred while exposed to the heavy forces of Hurricane Dean/Category 1.

Jules Craynock

Friday, September 14, 2007

Field Trip Completed

A team from AOML was at DBJM1 during the week of September 10th - 14th. The team consisted of Jules Craynock, Derek Manzello, Lew Gramer and Mike Jankulak.

Aerial work completed: the Vaisala Weather station mounting was adjusted by approximately nine degrees. The five solar panel cables were significantly shortened. The failing windbird / electronic compass combination was removed from the station and its aluminum mast put into storage. The station "brain" was removed and replaced; the logger firmware was upgraded on land and the internal barometer replaced. The barometer air tube was found to be filled with water at the outside end; this tube was removed. No damage of any kind (from Hurricane Dean) was observed above the surface. Four aluminum rungs were bent during station work but there are enough spare rungs to cover this loss.

Underwater work included adding 2-part epoxy to stabilize pin #8. Various maintenance tasks and biological survey activities were also performed (details to follow). The Groundtruth CT was connected during brain replacement to ensure that all connections were correct.

Future work: another trip will need to be scheduled to install the station "stabilization collar" and the replacement windbird and electronic compass.

-- Mike J+

Thursday, September 13, 2007

NO bleaching

I forgot to mention in my recent post that I observed little to no bleaching. There are some anecdotal reports of Agaricids starting to pale and/or bleach in deep water, but we were unable to observe this. Nevertheless, this sounds similar to the deep paling and partial bleaching of the plating Agaricids I saw in St. Croix two weeks ago. In short, nothing to write home about on the bleaching front.

FYI..in my previous post I mention the acronym YBD without defining it. YBD = Yellow Band Disease.

Cheerio,
Derek

Biological monitoring

Video transect surveys and urchin counts were performed today by Hugh Small, Camilo Trench, and myself. Fifteen video transects were run (16 m length transect line). All urchins within 0.5 m on either side of the transect line were identified and counted. A total of 20 urchin count transects were run (16 square meters per transect). Diadema antillarum is locally abundant near the ICON site and was the most abundant urchin surveyed. This is very encouraging as algal biomass is significantly reduced around the grazing area of individuals and where D. antillarum is absent, algal biomass is incredible (thats a scientific term).

I noted several large Acropora palmata pieces (up to 1 m in max. diameter) that had been broken and transported behind the reef crest during Hurricane Dean. Other than this, there were no indications of any negative hurricane impacts.

Nineteen colonies of Montastraea faveolata were tagged and photographed. Photos should be re-taken in 3 months to assess mortality and recovery. YBD appears to only be affecting a small proportion of M. faveolata colonies, but diseased colonies appear to have a clumped distribution. Colonies were tagged because YBD seems to be occurring throughout the Caribbean at the present time. I have received confirmed reports of YBD in the Dominican Republic and U.S. Virgin Islands to date. YBD appears to have been more abundant at the St. Croix site.

I would like to extend my utmost gratitude for the field support provided by Hugh and Camilo and to our captain Anthony.

Best regards,
Derek Manzello

Thursday, August 30, 2007

Station Cleaning

Full station cleaning carried out on 30th Aug.

Calibration unit secured to mast at depth of shallow CTD between 1100 – 1700 hrs local time.

Cleaning took place between 1315 – 1440 hrs local time.

Screens on CTD units changed

Algal growth not heavy

Thursday, August 23, 2007

Post-hurricane Inspection

Post hurricane inspection of station carried out 23 Aug 07. All intact except for outer North & North-Eastern stanchions. Cement holding pin for north-eastern stanchion badly cracked and stanchion moveable (shaken side to side) by hand but still holding and working to stabilise unit.
Cement holding pin for northern stanchion cracked but pin still secure and not moveable by hand
Good work by Jules et al

Monday, August 20, 2007

Post-Hurricane Dean Entry


The station withstood Hurricane Dean during the night and early morning of August 19th/20th. The maximum averaged hourly gusts measured at the station were 67.7 knots (77.9 mph), while maximum averaged hourly sustained winds were 48.5 knots (55.8 mph). The lowest barometer reading (hourly average) was about 997 millibars. Individual readings throughout each hour were no doubt higher and lower than these hourly averages. The station with calibrated instruments was just installed in June, so the data should be pretty accurate. (Click on graph above for enlarged view; times are in UTC.)

So far as I am aware, these meteorological and oceanographic data are the most comprehensive (i.e., to include sea temperature, salinity, light, pCO2, winds, etc.) for a coral reef area before, during and after a hurricane. A biological survey will be conducted soon.

We hope for the best for all of Jamaica's citizens and their country. We have not heard from Peter Gayle of Discovery Bay Marine Laboratory yet.

Jim

Friday, August 17, 2007

Maintenance Updates

There have been two cleanings since the first one, but Peter Gayle has not been able to enter complete records yet. They are preparing for Hurricane Dean.

Thursday, August 16, 2007

Basic Station Cleaning

Basic station cleaning carried out on 16 Aug 07.

Thursday, July 19, 2007

Full Station Cleaning

Full station cleaning carried out on 19th July 07.

Calibration unit secured to mast at depth of shallow CTD between 0945 and 1700 hrs local time.

Cleaning took place between 1355 – 1505 hrs local time.

Juvenile file fish trapped inside screen for CO2 sensor which cannot be removed completely for proper cleaning because of instrument / mounting hardware design.

Algal growth not heavy

Thursday, June 21, 2007

1st servicing

Operation : Basic Cleaning
Date : Wednesday, June 20, 2007
Time: 1102 - 1128
Observation: High sediment deposit on instrumments and supporting structures.
Very little algal growth but few patches of hydoids on lines/cables
Light Sensors: Cleaned sensor windows with chamois. 1102 - 1107 local time
Used soft toothbrush to remove sediment from the inner grooves and outside of cases.
The deep sensor appears to have an impression (approx. 2mm long and <0.5mm>
CTDs : Removed screens and brushed away sediment with soft toothbrush.
Cleaned outercase also. Slackened Tie wraps were tightened.
SAMI: Outer case and mesh cleaned and brushed
Suppoting stuctures : Cleaned with chamois and brushes.
Pylon, Spectra line, chains and cables were not serviced.

Labels:

Thursday, June 07, 2007

Station Operating!


By 10am local time, June 7, 2007, all the instruments were in place and the station was transmitting. The hourly in situ (only) data are available here, and the in situ data integrated with other data, plus a lot of other information on Discovery Bay Marine Lab (in background), West Fore Reef, the instruments, and the funding for this operation can be found here. Photo by Bernadette Charpentier.

Jim

Instrument Installation Team


From left to right, Chris Langdon (UM/RSMAS), Jim Hendee (NOAA/AOML), Mike Jankulak (UM/CIMAS), Nancy Ash (NOAA/AOML), and Anthony Downes (skipper, DBML).

(Photo by Bernadette Charpentier)

Congratulations!


Peter Gayle, Principal Scientific Officer of Discovery Bay Marine Laboratory (DBML), and Jim Hendee, Program Manger for ICON/CREWS, shake hands just after the completion of the installation of the CREWS station at West Fore Reef near DBML.

Photo by Bernadette Charpentier.

Instrument Layout



Layout of the instruments, just after completion. Click to enlarge.

Photo by Bernadette Charpentier.

Wednesday, June 06, 2007

Underwater Layout


A view from about 30' away shows Nancy Ash and Jim Hendee attending to preliminary wiring of the station. The eight chains are fastened to the bottom at radii of 30' away from the center to meet at stainless steel hounds on the pylon just below the water level. June 6, 2007.

Jim

Photo by Bernadette Charpentier

SAMI pCO2 instrument installed


Early this morning Chris Langdon and Nancy Ash installed the SAMI pCO2 at about 5m depth. Click image to see larger.

Photo by Bernadette Charpentier.

Bottom of the Pylon

The bottom of the pylon sits on a round 2" stainless steel trailer-hitch welded to a stainless steel plate fastened to the bottom, and is held down to the plate with extremely strong Spectra line. The weight of the chains up top also help to keep the station pinned down to the ball, and also serve as shock-absorbers during heavy seas and winds.

Note the zinc bar welded to the bottom plate, and the grounding plate positioned vertically on the pylon (wiring inside runs up to a lightening arrestor at the very top of the pylon). Click image to enlarge.

Photo by Bernadette Charpentier

Stringing Cables



The instruments are temporarily fastened to the stick on the first day so that Mike Jankulak, aloft on the pylon, can begin testing data throughput to the data logger.

Depth of the station is 20', top of the station is 18' above water.

Photo by Bernadette Charpentier.

Underwater Work


Nancy Ash and Jim Hendee work on preliminary attachment of instrument cables to the pylon, June 6, 2007, so that Mike Jankulak can test the wiring to the data logger aloft.

Jim

(Photo by Bernadette Charpentier)

Brain Surgery





Mike Jankulak attends to the wiring of the "brain" (i.e., the data logger), satellite transmitter, radio transceiver, GPS antenna, and various meteorological instruments.

Photo by Bernadette Charpentier

Progress



The "brain" (data logger, serial connections, etc.) was installed today, along with temporary placements of the instruments to get readings. We had the ever present ocean swell, with winds reaching 20 knots by noon, but mostly workable.

Image shown is Mike Jankulak, aloft, hoisting the RM Young Windbird from Chris Langdon.

Chris Langdon, Nancy Ash, Mike Jankulak and Jim Hendee